OpenStudy (anonymous):

Six pounds of cashews and eight pounds of peanuts cost $26.04. Four pounds of cashews and ten pounds of peanuts cost $21.59. Which system of equations can be used to determine the cost per pound of cashews, x, and the cost per pound of peanuts, y?

OpenStudy (anonymous):

6x+8y=26.04 4x+10y=21.59

OpenStudy (anonymous):

thank you

OpenStudy (anonymous):

An electronics store had a special sale on two different models of calculators. The more high-tech model sold for $33.04 and the older model sold for $16.88. All together, 29 calculators were sold. Total sales came out to $731.92. Which system of equations can be used to determine how many high-tech models were sold, x, and how many older models were sold, y?

OpenStudy (anonymous):

how far can u get?

OpenStudy (anonymous):

what?

OpenStudy (anonymous):

can't you write at least one of the equations?

OpenStudy (anonymous):

An electronics store had a special sale on two different models of calculators. The more high-tech model sold for $33.04 and the older model sold for $16.88. All together, 29 calculators were sold. Total sales came out to $731.92. Which system of equations can be used to determine how many high-tech models were sold, x, and how many older models were sold, y? A. 16.88y = 33.04x + 29 x + y = 731.92 B. 33.04x + 16.88y = 731.92 x - y = 29 C. 33.04x + 16.88y = 29 x + y = 731.92 D. 33.04x + 16.88y = 731.92 x + y = 29